1.Two point charges, Q= +8 μC and QB = -5 μC, are separated by a distance r = 10 cm. What is the magnitude of the electric force? The constant k = 8.988 x 109 Nm2/C


2.Two charged particles QP = +10 μC and Qq = +20 μC are separated by a distance r = 10 cm along the horizontal. What is the magnitude of the electrostatic force?


1
Expert's answer
2022-03-08T08:56:59-0500

The Coulomb's law says


"F=k\frac{q_1q_2}{r^2}"


1.

"F_1=9*10^9*\frac{8*10^{-6}*5*10^{-6}}{(0.1)^2}=36\:\rm N"

2.

"F_2=9*10^9*\frac{10*10^{-6}*20*10^{-6}}{(0.1)^2}=180\:\rm N"